OpenStudy (anonymous):

Solve system: y = x^2 - 3x y = 2x - 6

OpenStudy (anonymous):

set them equals to each other x^2-3x = 2x -6

OpenStudy (anonymous):

How do I solve?

OpenStudy (anonymous):

add 3x to both side

OpenStudy (anonymous):

x^2-5x+6=0 x^2-3x-2x+6=0 x(x-3)-2(x-3)=0 (x-2)(x-3)=0 So x=2,3

OpenStudy (anonymous):

What about y....?